OpenStudy (anonymous):

a golf ball dropped into a hard surface rebounds excally 1/3 of the height from which it falls, how far will it rebound after hitting the surface for the third time if it is dropped from a height of 18m?

OpenStudy (jack1):

=18m x 1/3 x 1/3 x 1/3 = 18 x 1/27 = 18/27